A "120-watt" lightbulb uses an average power of 120 W when connected to an rms voltage of 120 V.

What is the resistance of the lightbulb?

What is the maximum current in the bulb?

What is the maximum power used by the bulb at any given instant of time?
Answer #1


power P=120 W

voltage V_rms=120 V

Vmax=V_rms*sqrt(2)

=120*1.414

=169.7

a)

P=V_rms^2/R

120=120^2/R

=====>

resistance R=120 ohm


b)

current i_rms=V_rms/R

=120/120

=1 A

maximum current, imax=i_rms*sqrt(2)

=1.414A

c)

Pmax=Vmax*imax

=169.7*1.414

=240 W